cot54/tan36+tan20/cot70

Grade:8

1 Answers

Vedant Tiwari
8 years ago
Using the concept of tan(90-X)=cotX and Similarly cot(90-X)=tanX
So,
cot54 is equal to tan36 where X=54; Similarly tan20 is equal to cot70 where X=20
From the above results:
=tan36/tan36 + cot70/cot70
=1 + 1
=2